### v3

# API Pack: A Massive Multi-Programming Language Dataset for API Call Generation

Abstract
We introduce API Pack, a massive multi-programming language dataset containing more than 1 million instruction-API call pairs to improve the API call generation capabilities of large language models. By fine-tuning CodeLlama-13B on 20,000 Python instances from API Pack, we achieved around 10% and 5% higher accuracy compared to GPT-3.5 and GPT-4, respectively, in generating unseen API calls. Fine-tuning on API Pack enables cross-programming language generalization by leveraging a large amount of data in one language and small amounts of data from other languages. Scaling the training data to 1 million instances further improves the model’s generalization to new APIs not encountered during training. We open-source the API Pack dataset, trained models, and associated source code at https://github.com/zguo0525/API-Pack to facilitate further research.
1 Introduction
Large language models (LLMs) have shown promise in assisting with software engineering tasks [1, 2, 3, 4, 5, 6, 7, 8, 9], with a primary focus on code generation [10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26]. Our work builds on these advances but targets at the time-consuming task developers often face in finding application programming interface (API) call code examples. While finding an appropriate API is relatively easy with search, the challenge lies in generating the specific API call code. This involves understanding the API’s parameters, data types, and required operations, which can be complex and vary significantly across different APIs. Currently, developers typically seek out examples in documentations or API hubs [27], sifting through lengthy pages to locate relevant information [28]. This process is recognized as cumbersome and inefficient, as it requires developers to manually parse and adapt the example code to their specific use case. Our research aims to transform this workflow by using LLMs to identify appropriate API endpoints and generate corresponding API calls based on natural language instructions, thus automating the process of understanding and adapting API call code examples.
To achieve this research goal, we created API Pack, a dataset designed to improve the API call generation capabilities of LLMs. This massive multi-programming language dataset, containing over 1 million instances across 10 programming languages, is by far the largest open-source instruction dataset (see Table 1) for API call generation and API call intent detection. We evaluated the effectiveness of API Pack in improving API call generation through various experiments (see Section 4), and find that CodeLlama-13B, fine-tuned with only 20,000 Python instances from API Pack, surpasses the performance of GPT-3.5 and GPT-4 in generating API calls for entirely new APIs.
API Pack distinguishes itself from prior works [29, 30, 31] in two key aspects: multilinguality and scale. API Pack includes API calls across 10 different programming languages (see Table 1), allowing for a unique assessment of cross-lingual skill transfer - the extent to which improvements in one language apply to others. This is crucial for the practical application of LLMs in API call generation but has been underexplored in previous studies. Furthermore, with over a million instances spanning a wider range of real-world APIs and use cases compared to its predecessors, API Pack enables a more rigorous evaluation of generalization capabilities by controlling the volume of training data.
We summarize three key findings from our experiments:
Table 1: A comparison of API Pack with other instruction datasets for API intent detection and/or API call code generation. The upper section of the table reports the features that each dataset covers, and the bottom section reports the data statistics available (/ means unavailable).
Feature API Pack (this work) APIBench (Gorilla) ToolBench ToolBench (ToolLLM) API Bank ToolAlpaca ToolFormer
API call intent detection?
API call code generation?
Multi-lingual API calls?
Multi-API call scenario?
Data generation method: custom self-instruct self-instruct custom custom custom custom
# of Sources 4 3 8 1 53 / 5
# of APIs / Tools 11,213 1,645 8 16,464 53 400 5
# of API calls 1,128,599 16,450 / 37,204 568 3,938 9,400
# of Instances 1,128,599 16,450 2,746 12,657 264 3,938 22,453
  1. 1.
    Fine-tuning CodeLlama-13B on 20,000 Python instances from API Pack yields over 10% and 5% higher accuracy than GPT-3.5 and GPT-4, respectively, for unseen API calls.
  2. 2.
    Cross-lingual API call generation can be achieved by fine-tuning models on a large amount of data in one language plus small amounts of data from other languages.
  3. 3.
    Scaling the instruction data to 1 million instances improves generalization to new APIs, confirming the benefits of a larger dataset.
The paper is organized as follows: Section 2 reviews related works. Section 3 describes the construction of API Pack dataset. Section 4 outlines the experimental design for fine-tuning models on API Pack and evaluating their performance. The key findings are presented in Section 5. We conclude the paper in Section 6, and address the limitations in Section 7. API Pack dataset is licensed under the Creative Commons Attribution 4.0 International (CC BY 4.0) License. The associated source code is licensed under an MIT License.
2 Related Works
2.1 Methods to Generate Instruction Data with LLMs
Manually creating instruction data is labor-intensive [32]. To automate this process, researchers have proposed using LLMs, despite the risk of generating low-quality instances. Two popular methods are Self-Instruct [33] and Evol-Instruct [32], and both methods involve instance generation and filtering. Self-Instruct mutates seed instances using in-context examples and filters them based on ROUGE-L similarity and other heuristics. Evol-Instruct generates instances via targeted prompting and filters them using predefined heuristics. Quality checks are crucial when fine-tuning models with synthetic data. Expert annotators can determine instance correctness [33], or strong LLMs like ChatGPT can score or classify instances [34] based on heuristics derived from a small sample. This approach has also been used to assess instruction complexity [35, 36].
2.2 LLMs for API Call Code Generation and Intent Detection
LLMs for code generation primarily focus on general coding tasks, evaluated by benchmarks like MBPP [37], HumanEval [38], and its variants [39, 40, 41]. Recent research explores LLMs and APIs in two areas: API call intent detection and API call code generation. API call intent detection identifies the appropriate API endpoint for a natural language task. LLMs for this purpose [31, 42, 43, 44, 45] work in hybrid architectures, where the LLM finds the API endpoint(s) and other components generate the code. These studies explore single and multi-API intent scenarios, with few achieving good results in the latter [31]. Our work focuses on creating LLMs that generate code to invoke API functionality, which is less explored than API call intent detection. Gorilla [30] generates API calls for loading pre-trained machine learning models from known model hubs. ToolBench [29] benchmarks tool manipulation capabilities of open-source LLMs. Another line of work explores function calling in language models [46, 47, 48], including ChatGPT. These models require pre-defined functions before execution, differing from our scenario where developers rely on the language model to guide them to the appropriate function or API based on natural language input.
3 API Pack
API Pack is an instruction dataset with more than one million instances. Each instance contains an input-output pair plus additional information about the API and respective endpoints. The inputs are instructions for finding an API call to solve a coding task, including a task description in software engineering language and the name of the API to be used. The outputs are API call examples, specifically HTTP request code snippets curated from OpenAPI specification (OAS) files. The data in API Pack is sourced from four hubs that store OAS files: RapidAPI [^1] https://rapidapi.com/categories, APIGurus [^2] https://apis.guru/, Swaggerhub [^3] https://app.swaggerhub.com/search, and IBM’s public API Hub [^4] https://developer.ibm.com/apis/. Table 2 summarizes the total number of APIs, unique endpoints, and total instances (in different programming languages) contained within API Pack.
Table 2: Final count of data curated per source, an instance has one API call.
Source APIs Unique Endpoints Total Instances
IBM API Hub 73 2,884 17,206
APIs Gurus 1,980 37,097 495,533
Swaggerhub 5,045 26,747 345,765
RapidAPI 4,115 21,525 270,095
Total 11,213 88,253 1,128,599
The construction of API Pack involves four main stages: data pre-processing (Section 3.1), API Database (DB) creation (Section 3.2), instruction generation (Section 3.3), and data validation (Section 3.4). Figure 1 illustrates the overall pipeline.
Figure 1: Dataset curation pipeline.
3.1 Data Pre-processing
The data pre-processing stage involves two main steps. First, we filter out OAS files with non-English data to maintain language consistency. We also remove OAS files with zero endpoints, as they lack the necessary information to generate meaningful API call examples or instructions. Next, we extract relevant information from the remaining OAS files. At the endpoint level, we collect the name, functionality, description, method, and path. At the API level, we gather the API name, description, and provider. To further ensure data quality, we apply a second filter to remove instances with missing crucial data for generating API calls (e.g., method, path, or endpoint name) and instructions (e.g., functionality, description, or endpoint name). This step guarantees that all instances in the dataset contain the necessary information for generating useful API call examples and instructions.
3.2 Create API Databases
We build an API database (DB) using the pre-processed data, with each instance containing endpoint information, API details, an API call example in a specific programming language, and a language identifier. We use OpenAPI Snippet [^5] https://www.npmjs.com/package/openapi-snippet to generate API calls in 10 programming languages (cURL, libcurl, Java, Node.js, Python, Go, Ruby, PHP, Swift, JavaScript) for endpoints from RapidAPI, APIGurus, and OpenAPI, while directly extracting API calls from OAS files for IBM API Hub. Appendix 9.5 provides details on the API DB instance structure and the programming language diversity for each source.
3.3 Instructions Generation
The instruction generation process consists of two main steps: creating high-quality instruction examples and generating instruction candidates. This two-step process is essential to ensure the generation of precise, clear, and useful instructions.
First, we create high-quality instruction examples for each API DB file. We randomly select three endpoints from each API and use their information (e.g., functionality, description, endpoint_name, path) along with the respective API name to fill in a list of predefined instruction templates. This process yields three instruction examples per API DB file. For the API Gurus and IBM API Hub sources, three authors manually review and refine the generated instruction examples by correcting grammatical errors, removing unnecessary information, and ensuring that the API name is present in all examples. During this manual review, we identify common error patterns in the generated instructions. To streamline the refinement process for the Swaggerhub and RapidAPI sources, we replace human labor with an LLM (Mistral-7B-Instruct-v0.2666https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.2). We create a prompt (Prompt 1 in Appendix 9.8) based on the common error patterns identified during the manual review of the API Gurus and IBM API Hub instruction examples. This prompt is used to guide the LLM in refining the instruction examples for the Swaggerhub and RapidAPI sources.
Next, we use the high-quality instruction examples to generate instruction candidates for each instance in the API DBs. We prompt an LLM (Mistral-7B-Instruct-v0.2) with the endpoint’s information as input and provide the high-quality instruction examples as in-context examples (Prompt 2 in Appendix 9.8). The LLM generates five instruction candidates for each instance in the API DBs. Figure 2 shows an example of a generated instruction candidate, while Appendix 9.7 presents all five candidates generated for the same instance.
{minted}
[fontsize=, breaklines, breakanywhere] bash "instruction_candidates": [ "idx": 1, "candidate": "I’d like to confirm if 3D Secure is supported for a specific card brand or BIN using the Adyen BinLookup API. For 3D Secure 2, it will also provide device fingerprinting keys.", "input_tokens_mean": -0.5497539341557909, , … ], "api_call": "curl –request POST  –url https://paltest.adyen.com/pal/servlet/BinLookup/v40/get3dsAvailability  –header ’Authorization: Basic REPLACE_BASIC_AUTH’  –header ’content-type: application/json’  –data ’"additionalData":"property1":"string","property2":"string", "brands":["string"], "cardNumber":"string", "merchantAccount":"string", "recurringDetailReference":"string", "shopperReference":"string"’" …
Figure 2: Fragment of an API Pack instance in cURL.
3.4 Data Validation
The data validation process involves three main steps: 1) verifying the validity of API calls as HTTP request examples in a given programming language, 2) assessing the quality of generated instructions, and 3) selecting the highest-quality instruction for fine-tuning.
To verify the validity of API calls, we first compare the API call’s content with the instance data to ensure the correctness of the endpoint_name and the HTTP method (e.g., get, post, put, delete, custom method name). We then use regular expressions to validate the URL format, considering that placeholder strings may exist for the URL domain, path parameters, or query parameters. Additionally, we check that the programming language keywords in the API call match the language string ID assigned to each instance.
To validate the generated instructions for each API call, we randomly select a sample of 121 instances, each containing one API call and five instructions. These instances have unique functionality, description, and endpoint_name. One of the authors manually labels the 605 instructions (121 instances x 5 instructions) as good or bad. The analysis of the instructions labeled as bad reveals three common characteristics: 1) containing multiple instructions instead of a single one, b) including unnecessary text before or after the main instruction, and c) failing to accurately use the correct API name. Based on these characteristics, we create three prompts with a fixed set of in-context examples to automatically label instructions. To select the best prompt, we use an LLM (Mistral-7B-Instruct-v0.2) to label the 605 manually classified instructions and compare the results with human-labeled results. We choose the prompt with the best performance (Prompt 4 in Appendix 9.8) and instruct Mistral-7B-Instruct-v0.2 to classify all instructions in our dataset as good or bad. Instances with fewer than two good instructions are removed from the dataset.
To select the best candidate from the subset of good instructions, we calculate the likelihood of an LLM re-generating the input text used to generate the instruction. We prompt Mistral-7B-Instruct-v0.2 with each instruction candidate (Prompt 3 in Appendix 9.8) and obtain the log probability of each token for the re-generated input text. We calculate the mean of these log probabilities (input_tokens_mean). Appendix 9.7 shows five instruction candidates and their respective input_tokens_mean. We select instructions labeled as good with the best input_token_mean for fine-tuning. Our final dataset contains 1,128,599 instances, each with a valid API call example and at least two high-quality instructions. Appendix 9.1 shows the data instances filtered out at each stage of the pipeline.
4 Experiments and Evaluation Framework
To optimize the instruction-following capabilities of the language models, we post-processed API Pack into two instruction-tuning templates: zero-shot and three-shot. The zero-shot template targets scenarios where the output is expected to be a straightforward inference from the given input. The three-shot template emphasizes the model’s ability to learn and generate output with in-context learning. The three-shot template is available in Appendix 5.
4.1 Experimental Settings
A. Selecting the baseline: Our first experimental setting serves the purpose of selecting a base model for the rest of our experiments. We fine-tune Mistral 7b [49], CodeLlama 7b and 13b, as well as Llama 2 13b [50] on a subset of API Pack (20,000 instances in Python programming language). We evaluate the performance of each resulting model fine-tuned on the 20k subset and use the base model of the best-performing fine-tuned model in the rest of our experiments.
B. Inference with retrieval: Our second experiment aims to understand the influence of retrieval augmentation on model generalization. We evaluate the models under four distinct prompt settings during test time:
  • 0-shot: No API examples provided for the model.
  • 3-shot random: Three randomly selected API examples.
  • 3-shot retrieved: Three retrieved relevant API examples.
  • 3-shot retrieved & re-ranked: Five retrieved API examples, with three selected using a re-ranker model.
Note that these prompt settings are for testing/inference and differ from the instruction-tuning templates used in fine-tuning. We use bge-large-en-v1.5 [51] as the embedding model for retrieval and beg-reranker-large [52] for re-ranking. Appendix 9.4 illustrates the inference pipelines (0-shot, 3-shot) used to evaluate the models’ performance.
C. Cross-lingual generalization: To test the model’s ability to generalize to new programming languages, we supplement a cURL dataset of 100,000 instances with 1,000 instances from each of the nine additional languages in API Pack: Go, Java, JavaScript, libcurl, Node.js, PHP, Python, Ruby, and Swift. The goal is to determine if a model can generalize to new languages without requiring a large amount of multi-programming language data.
D. Scaling experiment: We conduct a scaling experiment to investigate whether more API data improves a model’s generalization ability for unseen API. We fine-tune models on progressively larger API datasets with unique API calls: 10k, 20k, 40k, 80k, 100k, and 1 million instances. Our hypothesis is that exposure to a greater scale and diversity of APIs during fine-tuning will improve the model’s ability to generalize to new, unseen APIs.
4.2 Evaluations
To measure the generalization capabilities enabled by API Pack, we establish a comprehensive evaluation framework spanning three levels of complexity for API call generation:
  • Level 1: Seen APIs and endpoints. This level assesses generalization to new instructions.
  • Level 2: Seen APIs and new endpoints. This level tests generalization to new endpoints of known APIs.
  • Level 3: Unseen APIs and endpoints. This level validates performance on entirely new APIs.
The process of data splitting is detailed in Appendix 9.2. Training hyperparameters are in Appendix 9.3. Endpoint and API call accuracy at each level is measured by SequenceMatcher, which identifies the longest matching subsequences while excluding insignificant elements. These elements refer to specific parts of the API requests that do not affect the overall functionality or purpose of the request, such as whitespace and formatting differences (e.g., variations in spaces, tabs, or newlines within the API requests) and specific naming conventions or variable names (i.e., the focus is on the structure and parameters of the API requests rather than the specific names used for variables or functions, as long as they serve the same purpose). A heuristic threshold of 0.9 is applied to compare the generated output with the ground truth Endpoint and API call, where the developers are expected to complete the API call based on the model response above this threshold.
5 Results
Table 3: Evaluation for models fine-tuned with 20k Python API dataset. A comparison with (non-fine-tuned) GPT-3.5 and GPT 4 performance for level 3 is also included.
Model Fine-tuning Testing Evaluation Accuracy (%)
template Level 1 Level 2 Level 3
Intent API Call Intent API Call Intent API Call
Mistral-7b zero-shot 0-shot 17.2 10.9 14.1 11.4 14.3 11.2
3-shot (retre) 42.0 29.7 35.4 28.7 39.1 29.1
three-shot 0-shot 40.5 28.5 24.0 18.3 15.2 12.1
3-shot (retre) 64.1 55.4 49.1 42.8 50.8 42.5
CodeLlama-7b zero-shot 0-shot 8.1 6.1 10.0 7.0 11.0 7.8
3-shot (retre) 52.6 42.6 43.6 35.9 50.2 40.1
three-shot 0-shot 12.1 9.3 13.7 10.2 16.8 13.0
3-shot (retre) 60.6 52.7 54.1 47.3 55.9 49.1
Llama-2-13b zero-shot 0-shot 9.4 6.2 11.6 9.0 10.9 8.4
3-shot (retre) 44.5 33.9 45.4 35.6 46.7 39.1
three-shot 0-shot 15.7 10.2 14.0 11.2 11.7 9.6
3-shot (retre) 59.5 51.5 50.8 44.3 52.7 44.2
CodeLlama-13b zero-shot 0-shot 9.8 6.8 10.8 8.1 12.1 8.5
3-shot (retre) 55.6 44.4 50.6 43.3 52.3 44.1
three-shot 0-shot 14.4 10.3 15.9 13.3 14.2 8.9
3-shot (retre) 63.5 55.5 56.8 51.4 56.1 49.5
gpt-3.5-1106 none 0-shot - - - - 1.0 0.7
none 3-shot (retre) - - - - 47.2 39.5
gpt-4-1106 none 0-shot - - - - 0.2 0.1
none 3-shot (retre) - - - - 53.5 44.3
5.1 Fine-tuned CodeLlama Excels in API call Generation
Table 3 presents the evaluation results for the four models fine-tuned with 20,000 Python instances from API Pack. The fine-tuned CodeLlama-13b model demonstrates superior performance in API call generation for the three-shot retrieved setting, achieving the highest API call accuracy across all evaluation levels. Models fine-tuned with the three-shot template consistently outperform those fine-tuned with the zero-shot template, showing the importance of using the three-shot template to improve the model’s in-context learning abilities. Furthermore, prompting with 3-shot (retre) at testing time leads to substantial improvements compared to 0-shot, a trend observed across all models and levels. This finding highlights the significance of providing relevant examples to improve the models’ accuracy in generating API calls. Notably, the fine-tuned CodeLlama-13b model with API Pack surpasses the performance of GPT-3.5 and GPT-4 models (not fine-tuned) for Level 3, both 0-shot and 3-shot (retre).
Table 4: Performance comparison for different retrieval methods in three-shot prompting.
Model Testing Evaluation Accuracy (%)
Level 1 Level 2 Level 3
Endpoint API Call Endpoint API Call Endpoint API Call
Mistral-7b 3-shot (rand) 54.5 41.8 48.2 41.2 45.2 37.0
3-shot (retre) 64.1 55.4 49.1 42.8 50.8 42.5
3-shot (retre & rerank) 63.0 53.6 49.0 42.2 51.5 43.9
CodeLlama-13b 3-shot (rand) 49.2 38.6 49.8 43.6 50.0 41.4
3-shot (retre) 63.5 55.5 56.8 51.4 56.1 49.5
3-shot (retre & rerank) 61.0 52.9 55.1 49.2 55.9 49.3
5.2 Retrieval Augmentation Improves API Call Generation
Table 4 illustrates the impact of different retrieval methods on three-shot API call generation, highlighting the significance of retrieval augmentation in improving the models’ performance. The results of fine-tuning Mistral-7b and CodeLlama-13b models with 20,000 Python instances from API Pack using the three-shot template demonstrate that three-shot retrieval is the most effective approach.
5.3 Cross-lingual Generalizations in API Call Performance
Figure 3: Cross-lingual API call performance for different models in cURL, Python, and Java.
Figure 4: Three-shot performance in ten languages.
Figure 3 compares three fine-tuning approaches for cross-lingual API call performance: a model fine-tuned exclusively on 100,000 instances of ‘cURL’ data, or cURL Model, three expert models, each fine-tuned on 100,000 samples of ‘cURL’, ‘Python’, and ‘Java’ data separately, and a mixture model fine-tuned on 100,000 instances of ‘cURL’ data with additional samples of 1,000 instances each for nine different languages.
The graph shows that generalizing to new programming languages without fine-tuning is challenging in zero-shot scenarios. However, the models perform better than expected in three-shot settings, indicating some in-context learning adaptability. Mixed-language fine-tuning improves performance in both zero-shot and three-shot scenarios, suggesting that even a small amount of fine-tuning data in various languages can improve the model’s effectiveness in cross-lingual API call tasks. These findings demonstrate that cross-lingual API call generation can be achieved by fine-tuning models on a large amount of data in one language plus small amounts of data from other languages, without requiring an extensive dataset for each target programming language. Figure 4 illustrates the models’ adaptability across ten programming languages, with the Mixture Model performing on par with Expert Models in three-shot testing, highlighting their potential in multi-lingual programming applications.
5.4 Scaling Instruction Dataset Helps Generalization
Figure 5: Scaling instruction dataset on CodeLlama-13b with zero-shot and three-shot retre evaluations. The x-axis represents the size of fine-tuning data from API Pack. The y-axis is Endpoint or API call accuracy.
Fig. 5 clearly demonstrates an upward trend in zero-shot performance that correlates with the fine-tuning dataset size (20k, 40k, 80k, 100k, and 1 million instances from API Pack). This improvement highlights the advantage of larger datasets in providing diverse examples, which is crucial in zero-shot prompting where the model relies solely on its pre-existing knowledge. For three-shot prompting, the graph also shows improvements as the dataset size increases, although less significantly than in the zero-shot case, particular for level 3 cases. This suggests that while additional fine-tuning data is beneficial, model pretraining plays a more critical role in in-context learning ability.
5.5 Additional Ablation Experiments
Table 5: Performance comparison for ToolBench and API Pack in different shot settings.
Train Testing Evaluation Accuracy
Level 1 Endpoint Level 1 API Call Level 2 Endpoint Level 2 API Call Level 3 Endpoint Level 3 API Call
ToolBench 0-shot 5.7 5.7 8.1 8.0 7.3 7.0
3-shot 44.5 37.8 40.7 36.4 43.7 38.1
API Pack 0-shot 14.4 10.3 15.9 13.3 14.2 8.9
3-shot 63.5 55.5 56.8 51.4 56.1 49.1
To verify API Pack’s quality, we evaluated a fine-tuned CodeLlama-13b model on 20k instances of ToolBench dataset, which only contains APIs from RapidAPI. Results in Table 5 show that using APIs from a single source like in TollBench decreases model performance in both 0-shot and 3-shot settings compared to API Pack, which curates data from four different sources.
Table 6: Performance comparison for none-filtered and with-filtered datasets in different shot settings.
Train Testing Evaluation Accuracy
Level 1 Endpoint Level 1 API Call Level 2 Endpoint Level 2 API Call Level 3 Endpoint Level 3 API Call
none-filtered 0-shot 10.3 8.3 12.8 10.3 12.2 8.4
3-shot 62.4 54.9 54.9 48.7 55.7 49.5
with-filtered 0-shot 14.4 10.3 15.9 13.3 14.2 8.9
3-shot 63.5 55.5 56.8 51.4 56.1 49.1
Furthermore, to demonstrate the effectiveness of our data filtering pipeline, we replaced 45% of the filtered instructions with non-filtered ones and fine-tuned CodeLlama-13b with 20k instances of each. The results in Table 6 confirm that the 0-shot and 3-shot performances indeed drop with the introduction of non-filtered instructions.
6 Conclusion
In this paper, we introduce API Pack, a massive multi-programming language dataset with over 1 million instruction-API call instances. Using this dataset, we investigate if models can generalize to new languages without extensive multi-language data and whether exposure to diverse APIs during fine-tuning improves generalization. Our results show that cross-lingual API call is achievable by training on a large amount of single-language data supplemented with small amounts of data from other languages. Furthermore, increasing data volume improves generalization. These findings highlight the potential of large-scale, diverse datasets in enabling LLMs to effectively identify appropriate API endpoints and generate corresponding API calls based on natural language instructions.
7 Limitations
API Pack is not designed for multi-API call scenarios, so models fine-tuned with this dataset may struggle with tasks involving multiple, interdependent API calls. This limitation restricts API Pack’s effectiveness in complex, real-world software development tasks that often require multi-API integrations. Future work should focus on creating datasets and models that can effectively handle multi-API call scenarios, enabling the generation of more sophisticated and practical code solutions.
8 Broader Impact Statement
Our research has a broader impact on Software Productivity as it aims to accelerate software development workflows by automating routine coding tasks. Alongside promising productivity gains, the integration of advanced LLMs into software development workflows raises sociotechnical concerns (e.g., job displacement, developers’ continuous oversight to ensure accuracy). Therefore, a responsible innovation lens assessing and responding to potential harms remains vital.
Acknowledgments and Disclosure of Funding
This work was supported by MIT-IBM Watson AI Lab and IBM Research.

### v4

@@
 We introduce API Pack, a massive multi-programming language dataset containing more than 1 million instruction-API call pairs to improve the API call generation capabilities of large language models.
-By fine-tuning CodeLlama-13B on 20,000 Python instances from API Pack, we achieved around 10% and 5% higher accuracy compared to GPT-3.5 and GPT-4, respectively, in generating unseen API calls.
-Fine-tuning on API Pack enables cross-programming language generalization by leveraging a large amount of data in one language and small amounts of data from other languages.
-Scaling the training data to 1 million instances further improves the model’s generalization to new APIs not encountered during training.
-We open-source the API Pack dataset, trained models, and associated source code at https://github.com/zguo0525/API-Pack to facilitate further research.
+By fine-tuning CodeLlama-13B on 20,000 Python instances from API Pack, we enable it to outperform GPT-3.5 and GPT-4 in generating unseen API calls.
+Fine-tuning on API Pack also facilitates cross-programming language generalization by leveraging a large amount of data in one language and small amounts of data from other languages.
+Scaling the training data to 1 million instances further improves the model’s ability to generalize to new APIs not used in training.
+To facilitate further research, we open-source the API Pack dataset, trained model, and associated source code at https://github.com/zguo0525/API-Pack.
 1 Introduction
 Large language models (LLMs) have shown promise in assisting with software engineering tasks [1, 2, 3, 4, 5, 6, 7, 8, 9], with a primary focus on code generation [10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26].
-Our work builds on these advances but targets at the time-consuming task developers often face in finding application programming interface (API) call code examples.
+Our work builds on these advances but targets a time-consuming task developers often face in finding application programming interface (API) call code examples.
 While finding an appropriate API is relatively easy with search, the challenge lies in generating the specific API call code.
@@
 Currently, developers typically seek out examples in documentations or API hubs [27], sifting through lengthy pages to locate relevant information [28].
-This process is recognized as cumbersome and inefficient, as it requires developers to manually parse and adapt the example code to their specific use case.
-Our research aims to transform this workflow by using LLMs to identify appropriate API endpoints and generate corresponding API calls based on natural language instructions, thus automating the process of understanding and adapting API call code examples.
-To achieve this research goal, we created API Pack, a dataset designed to improve the API call generation capabilities of LLMs.
-This massive multi-programming language dataset, containing over 1 million instances across 10 programming languages, is by far the largest open-source instruction dataset (see Table 1) for API call generation and API call intent detection.
-We evaluated the effectiveness of API Pack in improving API call generation through various experiments (see Section 4), and find that CodeLlama-13B, fine-tuned with only 20,000 Python instances from API Pack, surpasses the performance of GPT-3.5 and GPT-4 in generating API calls for entirely new APIs.
-API Pack distinguishes itself from prior works [29, 30, 31] in two key aspects: multilinguality and scale.
+This process which requires developers to manually look for code examples for their specific use case is often cumbersome and tedious.
+Our work aims to transform this workflow by using LLMs to identify appropriate API endpoints and generate corresponding API calls based on natural language instructions, thus automating the process of searching the appropriate API call code examples.
+To achieve this goal, we create API Pack, a dataset designed to improve the API call generation capabilities of LLMs.
+This multi-programming language dataset, containing over 1 million instances across 10 programming languages, is by far the largest open-source instruction dataset (see Table 1) for single API call generation and API call intent detection.
+API Pack distinguishes itself from prior works [29, 30, 31] in two key aspects: programming language diversity and dataset size.
 API Pack includes API calls across 10 different programming languages (see Table 1), allowing for a unique assessment of cross-lingual skill transfer - the extent to which improvements in one language apply to others.
-This is crucial for the practical application of LLMs in API call generation but has been underexplored in previous studies.
-Furthermore, with over a million instances spanning a wider range of real-world APIs and use cases compared to its predecessors, API Pack enables a more rigorous evaluation of generalization capabilities by controlling the volume of training data.
+This important aspect has been under explored in previous studies even though it is crucial for the practical application of LLMs to API call generation use cases.
+Furthermore, with more than a million instances that span a wider range of real-world APIs and use cases compared to its predecessors (see Table 1), API Pack enables a more rigorous evaluation of generalization capabilities by controlling the volume of training data.
 We summarize three key findings from our experiments:
@@
 API call code generation?
-Multi-lingual API calls?
+Multi-programming language? ✓(10) ✗(Python) ✓(Curl, Python)
 Multi-API call scenario?
-Data generation method: custom self-instruct self-instruct custom custom custom custom
+Data generation method custom self-instruct self-instruct custom custom custom custom
 # of Sources 4 3 8 1 53 / 5
@@
   2. 2.
-    Cross-lingual API call generation can be achieved by fine-tuning models on a large amount of data in one language plus small amounts of data from other languages.
+    Cross-lingual API call generation can be facilitated by fine-tuning models on a large amount of data in one language along with small amounts of data from other languages.
   3. 3.
-    Scaling the instruction data to 1 million instances improves generalization to new APIs, confirming the benefits of a larger dataset.
+    Scaling the instruction data from smaller instances to 1 million instances improves generalization to new APIs, validating the benefits of a larger dataset.
 The paper is organized as follows: Section 2 reviews related works.
@@
 The key findings are presented in Section 5.
-We conclude the paper in Section 6, and address the limitations in Section 7.
-API Pack dataset is licensed under the Creative Commons Attribution 4.0 International (CC BY 4.0) License.
-The associated source code is licensed under an MIT License.
+We conclude the paper in Section 6, and address limitations and future work plans in Section 7.
+We discuss ethical considerations related to API Pack’s use in Appendix 9.1.
+API Pack dataset is under the Creative Commons Attribution 4.0 International License.
+The associated source code is under an MIT License.
 2 Related Works
 2.1 Methods to Generate Instruction Data with LLMs
-Manually creating instruction data is labor-intensive [32].
-To automate this process, researchers have proposed using LLMs, despite the risk of generating low-quality instances.
-Two popular methods are Self-Instruct [33] and Evol-Instruct [32], and both methods involve instance generation and filtering.
+Manually creating instruction data is labor intensive [32].
+To automate this process, researchers have proposed the use of LLMs.
+Two popular methods to generate data via an LLM are Self-Instruct [33] and Evol-Instruct [32], both methods involve instance (instruction - response pairs) generation and filtering.
 Self-Instruct mutates seed instances using in-context examples and filters them based on ROUGE-L similarity and other heuristics.
 Evol-Instruct generates instances via targeted prompting and filters them using predefined heuristics.
-Quality checks are crucial when fine-tuning models with synthetic data.
-Expert annotators can determine instance correctness [33], or strong LLMs like ChatGPT can score or classify instances [34] based on heuristics derived from a small sample.
-This approach has also been used to assess instruction complexity [35, 36].
+Custom LLM-based pipelines to generate synthetic data, such as the one we followed to create API Pack, also exist.
+In the case of code datasets, the pipelines typically involve mining high-quality code and using an LLM to generate instructions for each code data sample (e.g., [30], [29]).
+Post-quality filtering is crucial when using LLM-based synthetic data for fine-tuning.
+Although human annotators can determine the correctness of the instance [33], this kind of revision takes time and is thus not reasonable for large datasets.
+Therefore, recent work has proposed the use of strong LLMs (i.e., GPT-4 and Mistral AI models families) to score or classify instances [34] based on heuristics and examples derived from a small sample of data reviewed by human annotators.
+A similar approach has also been used to assess instruction complexity [35, 36].
 2.2 LLMs for API Call Code Generation and Intent Detection
@@
 These studies explore single and multi-API intent scenarios, with few achieving good results in the latter [31].
-Our work focuses on creating LLMs that generate code to invoke API functionality, which is less explored than API call intent detection.
+Our work focuses on training LLMs that generate code to invoke API functionality, which is less explored than API call intent detection.
 Gorilla [30] generates API calls for loading pre-trained machine learning models from known model hubs.
@@
 Another line of work explores function calling in language models [46, 47, 48], including ChatGPT.
-These models require pre-defined functions before execution, differing from our scenario where developers rely on the language model to guide them to the appropriate function or API based on natural language input.
+These models require pre-defined functions before execution, differing from our scenario where developers rely on a language model to guide them to the appropriate function or API.
 3 API Pack
-API Pack is an instruction dataset with more than one million instances.
+API Pack is an instruction dataset with over one million instances.
 Each instance contains an input-output pair plus additional information about the API and respective endpoints.
-The inputs are instructions for finding an API call to solve a coding task, including a task description in software engineering language and the name of the API to be used.
+The inputs are instructions for finding an API call to solve a coding task, including a task description in software engineering language and the name of the API.
 The outputs are API call examples, specifically HTTP request code snippets curated from OpenAPI specification (OAS) files.
 The data in API Pack is sourced from four hubs that store OAS files: RapidAPI [^1] https://rapidapi.com/categories, APIGurus [^2] https://apis.guru/, Swaggerhub [^3] https://app.swaggerhub.com/search, and IBM’s public API Hub [^4] https://developer.ibm.com/apis/.
-Table 2 summarizes the total number of APIs, unique endpoints, and total instances (in different programming languages) contained within API Pack.
+To the best of our knowledge, these API Hubs were the only sources with publicly available OAS files at a large scale at the time we collected data.
+From these sources, we kept all the APIs that passed our filtering criteria (see Section 3.1).
+Table 2 summarizes the total number of APIs, unique endpoints, and total instances in API Pack.
 Table 2: Final count of data curated per source, an instance has one API call.
@@
 3.1 Data Pre-processing
-The data pre-processing stage involves two main steps.
-First, we filter out OAS files with non-English data to maintain language consistency.
-We also remove OAS files with zero endpoints, as they lack the necessary information to generate meaningful API call examples or instructions.
+This process involves two steps.
+First, as the presence of non-English metadata (e.g., API description, endpoint description) in the files was insignificant (less than 1%), we filter out OAS files with non-English data.
+We also remove OAS files with zero endpoints, as it is not possible to generate API calls for these APIs.
 Next, we extract relevant information from the remaining OAS files.
+We extracted information about the APIs, and about the endpoints that each API contains.
+At the API level, we gather the API name, description, and provider.
 At the endpoint level, we collect the name, functionality, description, method, and path.
-At the API level, we gather the API name, description, and provider.
 To further ensure data quality, we apply a second filter to remove instances with missing crucial data for generating API calls (e.g., method, path, or endpoint name) and instructions (e.g., functionality, description, or endpoint name).
@@
 3.2 Create API Databases
-We build an API database (DB) using the pre-processed data, with each instance containing endpoint information, API details, an API call example in a specific programming language, and a language identifier.
-We use OpenAPI Snippet [^5] https://www.npmjs.com/package/openapi-snippet to generate API calls in 10 programming languages (cURL, libcurl, Java, Node.js, Python, Go, Ruby, PHP, Swift, JavaScript) for endpoints from RapidAPI, APIGurus, and OpenAPI, while directly extracting API calls from OAS files for IBM API Hub.
+We build an API database (DB) using the pre-processed data, with each instance containing endpoint information, and API details.
+Then, we use OpenAPI Snippet [^5] https://www.npmjs.com/package/openapi-snippet library to generate API calls in 10 programming languages (cURL, libcurl, Java, Node.js, Python, Go, Ruby, PHP, Swift, JavaScript) for each API DB instance.
+We followed this process to create API calls with the data from RapidAPI, APIGurus, and OpenAPI.
+We directly extract API calls from OAS files for IBM API Hub as this source contained API calls in more than one programming language.
+Two factors influenced our selection of programming languages.
+First, the 10 programming languages included in API Pack appear in the StarCoder report of programming languages [23].
+Second, given that we used an open-source library (OpenAPI Snippet) to generate API calls, language programming selection is inherently limited to the languages that this library supports.
 Appendix 9.6 provides details on the API DB instance structure and the programming language diversity for each source.
 3.3 Instructions Generation
-The instruction generation process consists of two main steps: creating high-quality instruction examples and generating instruction candidates.
-This two-step process is essential to ensure the generation of precise, clear, and useful instructions.
-First, we create high-quality instruction examples for each API DB file.
-We randomly select three endpoints from each API and use their information (e.g., functionality, description, endpoint_name, path) along with the respective API name to fill in a list of predefined instruction templates.
+This process involves creating high-quality instruction examples, followed by generating instruction candidates based on these examples.
+First, we randomly select three endpoints from each API DB and use their information (e.g., functionality, description, endpoint_name, path) along with the respective API name to fill in a list of predefined instruction templates.
 This process yields three instruction examples per API DB file.
-For the API Gurus and IBM API Hub sources, three authors manually review and refine the generated instruction examples by correcting grammatical errors, removing unnecessary information, and ensuring that the API name is present in all examples.
-During this manual review, we identify common error patterns in the generated instructions.
-To streamline the refinement process for the Swaggerhub and RapidAPI sources, we replace human labor with an LLM (Mistral-7B-Instruct-v0.2666https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.2).
-We create a prompt (Prompt 1 in Appendix 9.8) based on the common error patterns identified during the manual review of the API Gurus and IBM API Hub instruction examples.
-This prompt is used to guide the LLM in refining the instruction examples for the Swaggerhub and RapidAPI sources.
-Next, we use the high-quality instruction examples to generate instruction candidates for each instance in the API DBs.
-We prompt an LLM (Mistral-7B-Instruct-v0.2) with the endpoint’s information as input and provide the high-quality instruction examples as in-context examples (Prompt 2 in Appendix 9.8).
-The LLM generates five instruction candidates for each instance in the API DBs.
-Figure 2 shows an example of a generated instruction candidate, while Appendix 9.7 presents all five candidates generated for the same instance.
-{minted}
-[fontsize=, breaklines, breakanywhere] bash "instruction_candidates": [ "idx": 1, "candidate": "I’d like to confirm if 3D Secure is supported for a specific card brand or BIN using the Adyen BinLookup API.
-For 3D Secure 2, it will also provide device fingerprinting keys.", "input_tokens_mean": -0.5497539341557909, , … ], "api_call": "curl –request POST  –url https://paltest.adyen.com/pal/servlet/BinLookup/v40/get3dsAvailability  –header ’Authorization: Basic REPLACE_BASIC_AUTH’  –header ’content-type: application/json’  –data ’"additionalData":"property1":"string","property2":"string", "brands":["string"], "cardNumber":"string", "merchantAccount":"string", "recurringDetailReference":"string", "shopperReference":"string"’" …
-Figure 2: Fragment of an API Pack instance in cURL.
+For the API Gurus and IBM API Hub sources, three authors manually refine the generated instruction examples by correcting grammatical errors, removing unnecessary information, and ensuring that the API name is present in all examples.
+To streamline the refinement process of examples for the Swaggerhub and RapidAPI sources, we replace human labor with an LLM (Mistral-7B-Instruct-v0.2666https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.2).
+We create a prompt to refine the instruction examples (Prompt 1 in Appendix 9.10) based on the heuristics identified during the manual review.
+We use this prompt to task the LLM with refining the instruction examples for the Swaggerhub and RapidAPI sources.
+Second, we use the high-quality instruction examples to generate five instruction candidates for each instance in the API DBs.
+We prompt an LLM (Mistral-7B-Instruct-v0.2) with the endpoint’s information as input and provide the high-quality instruction examples as in-context examples (see Prompt 2 in Appendix 9.10).
+Figure 7 in Appendix 9.8 shows an instance containing one instruction candidate along with its respective API call.
+Appendix 9.9 presents all five candidates generated for the same instance.
 3.4 Data Validation
@@
 To validate the generated instructions for each API call, we randomly select a sample of 121 instances, each containing one API call and five instructions.
-These instances have unique functionality, description, and endpoint_name.
-One of the authors manually labels the 605 instructions (121 instances x 5 instructions) as good or bad.
-The analysis of the instructions labeled as bad reveals three common characteristics: 1) containing multiple instructions instead of a single one, b) including unnecessary text before or after the main instruction, and c) failing to accurately use the correct API name.
-Based on these characteristics, we create three prompts with a fixed set of in-context examples to automatically label instructions.
-To select the best prompt, we use an LLM (Mistral-7B-Instruct-v0.2) to label the 605 manually classified instructions and compare the results with human-labeled results.
-We choose the prompt with the best performance (Prompt 4 in Appendix 9.8) and instruct Mistral-7B-Instruct-v0.2 to classify all instructions in our dataset as good or bad.
+These instances have unique functionality, description, and endpoint name.
+One of the authors manually annotates 605 instructions (121 instances x 5 instructions) as good or bad.
+The analysis of the instructions annotated as bad reveals three common characteristics: 1) containing multiple instructions instead of a single one, b) including unnecessary text before or after the main instruction, and c) failing to accurately use the correct API name.
+Based on these characteristics, we create three prompts with a fixed set of in-context examples to automatically annotate instructions via an LLM.
+To select the best prompt, we use an LLM (Mistral-7B-Instruct-v0.2) to annotate the 605 manually classified instructions and compare the results with human-annotated results.
+We choose the prompt with the best performance (Prompt 4 in Appendix 9.10) and instruct Mistral-7B-Instruct-v0.2 to classify all instructions in our dataset.
 Instances with fewer than two good instructions are removed from the dataset.
@@
 Appendix 9.9 shows five instruction candidates and their respective input_tokens_mean.
-We select instructions labeled as good with the best input_token_mean for fine-tuning.
+We select instructions annotated as good with the best input_token_mean for fine-tuning.
 Our final dataset contains 1,128,599 instances, each with a valid API call example and at least two high-quality instructions.
@@
 4 Experiments and Evaluation Framework
-To optimize the instruction-following capabilities of the language models, we post-processed API Pack into two instruction-tuning templates: zero-shot and three-shot.
-The zero-shot template targets scenarios where the output is expected to be a straightforward inference from the given input.
-The three-shot template emphasizes the model’s ability to learn and generate output with in-context learning.
-The three-shot template is available in Appendix 5.
+To optimize the instruction-following capabilities of the language models, we post-process API Pack into two instruction-tuning templates.
+The 0-shot-tuning template targets scenarios where the output is expected to be a straightforward inference from the given input.
+The 3-shot-tuning template, with random selections within the selected API, emphasizes the model’s ability to learn and generate output with in-context learning.
+The 3-shot-tuning template is available in Appendix 5.
 4.1 Experimental Settings
 A.
-Selecting the baseline: Our first experimental setting serves the purpose of selecting a base model for the rest of our experiments.
+Selecting a base model: Our first experimental study serves the purpose of selecting a base model for the rest of our experiments.
 We fine-tune Mistral 7b [49], CodeLlama 7b and 13b, as well as Llama 2 13b [50] on a subset of API Pack (20,000 instances in Python programming language).
-We evaluate the performance of each resulting model fine-tuned on the 20k subset and use the base model of the best-performing fine-tuned model in the rest of our experiments.
+We evaluate the performance of resulting models and select the best-performing base model.
 B.
@@
   • 3-shot retrieved: Three retrieved relevant API examples.
-  • 3-shot retrieved & re-ranked: Five retrieved API examples, with three selected using a re-ranker model.
-Note that these prompt settings are for testing/inference and differ from the instruction-tuning templates used in fine-tuning.
+  • 3-shot retrieved & re-ranked: Five retrieved API examples, with three selected using a re-ranker model in the final prompt.
+The 3-shot random inference uses the same prompt template as 3-shot-tuning.
 We use bge-large-en-v1.5 [51] as the embedding model for retrieval and beg-reranker-large [52] for re-ranking.
-Appendix 9.4 illustrates the inference pipelines (0-shot, 3-shot) used to evaluate the models’ performance.
+Appendix 9.5 illustrates the inference pipelines used to evaluate the models’ performance.
 C.
@@
 D.
-Scaling experiment: We conduct a scaling experiment to investigate whether more API data improves a model’s generalization ability for unseen API.
-We fine-tune models on progressively larger API datasets with unique API calls: 10k, 20k, 40k, 80k, 100k, and 1 million instances.
+Scaling experiment: We conduct a scaling experiment to investigate whether more API data improves a model’s generalization ability for unseen APIs.
+We fine-tune models on progressively larger API datasets with unique API calls on 10k, 20k, 40k, 80k, 100k, and 1 million instances.
 Our hypothesis is that exposure to a greater scale and diversity of APIs during fine-tuning will improve the model’s ability to generalize to new, unseen APIs.
@@
 A heuristic threshold of 0.9 is applied to compare the generated output with the ground truth Endpoint and API call, where the developers are expected to complete the API call based on the model response above this threshold.
-5 Results
-Table 3: Evaluation for models fine-tuned with 20k Python API dataset. A comparison with (non-fine-tuned) GPT-3.5 and GPT 4 performance for level 3 is also included.
+5 Experimental Results
+5.1 Fine-tuned CodeLlama Excels in API call Generation
+Table 3 presents the evaluation results for the four models fine-tuned with 20,000 Python instances from API Pack.
+The fine-tuned CodeLlama-13b model demonstrates superior performance in API call generation for the 3-shot retrieved setting, achieving the highest API call accuracy across all evaluation levels.
+Models fine-tuned with the 3-shot-tuning template consistently outperform those fine-tuned with the 0-shot-tuning template, showing the importance of using the 3-shot template to improve the model’s in-context learning abilities during training.
+Furthermore, prompting with 3-shot (retre) at testing time leads to substantial improvements compared to 0-shot, a trend observed across all models and levels.
+This finding highlights the significance of providing relevant examples to improve the models’ accuracy in generating API calls.
+Notably, the fine-tuned CodeLlama-13b model with API Pack surpasses the performance of GPT-3.5 and GPT-4 models (not fine-tuned) for Level 3, both 0-shot and 3-shot (retre).
+Table 3: Evaluation for models fine-tuned with 20k Python API dataset, including a comparison with (not-fine-tuned) GPT-3.5 and GPT 4. Note that 3-shot (retre) is short for 3-shot retrieved.
 Model Fine-tuning Testing Evaluation Accuracy (%)
@@
 Intent API Call Intent API Call Intent API Call
-Mistral-7b zero-shot 0-shot 17.2 10.9 14.1 11.4 14.3 11.2
+Mistral-7b 0-shot 0-shot 17.2 10.9 14.1 11.4 14.3 11.2
 3-shot (retre) 42.0 29.7 35.4 28.7 39.1 29.1
-three-shot 0-shot 40.5 28.5 24.0 18.3 15.2 12.1
+3-shot 0-shot 40.5 28.5 24.0 18.3 15.2 12.1
 3-shot (retre) 64.1 55.4 49.1 42.8 50.8 42.5
-CodeLlama-7b zero-shot 0-shot 8.1 6.1 10.0 7.0 11.0 7.8
+CodeLlama-7b 0-shot 0-shot 8.1 6.1 10.0 7.0 11.0 7.8
 3-shot (retre) 52.6 42.6 43.6 35.9 50.2 40.1
-three-shot 0-shot 12.1 9.3 13.7 10.2 16.8 13.0
+3-shot 0-shot 12.1 9.3 13.7 10.2 16.8 13.0
 3-shot (retre) 60.6 52.7 54.1 47.3 55.9 49.1
-Llama-2-13b zero-shot 0-shot 9.4 6.2 11.6 9.0 10.9 8.4
+Llama-2-13b 0-shot 0-shot 9.4 6.2 11.6 9.0 10.9 8.4
 3-shot (retre) 44.5 33.9 45.4 35.6 46.7 39.1
-three-shot 0-shot 15.7 10.2 14.0 11.2 11.7 9.6
+3-shot 0-shot 15.7 10.2 14.0 11.2 11.7 9.6
 3-shot (retre) 59.5 51.5 50.8 44.3 52.7 44.2
-CodeLlama-13b zero-shot 0-shot 9.8 6.8 10.8 8.1 12.1 8.5
+CodeLlama-13b 0-shot 0-shot 9.8 6.8 10.8 8.1 12.1 8.5
 3-shot (retre) 55.6 44.4 50.6 43.3 52.3 44.1
-three-shot 0-shot 14.4 10.3 15.9 13.3 14.2 8.9
+3-shot 0-shot 14.4 10.3 15.9 13.3 14.2 8.9
 3-shot (retre) 63.5 55.5 56.8 51.4 56.1 49.5
@@
 none 3-shot (retre) - - - - 53.5 44.3
-5.1 Fine-tuned CodeLlama Excels in API call Generation
-Table 3 presents the evaluation results for the four models fine-tuned with 20,000 Python instances from API Pack.
-The fine-tuned CodeLlama-13b model demonstrates superior performance in API call generation for the three-shot retrieved setting, achieving the highest API call accuracy across all evaluation levels.
-Models fine-tuned with the three-shot template consistently outperform those fine-tuned with the zero-shot template, showing the importance of using the three-shot template to improve the model’s in-context learning abilities.
-Furthermore, prompting with 3-shot (retre) at testing time leads to substantial improvements compared to 0-shot, a trend observed across all models and levels.
-This finding highlights the significance of providing relevant examples to improve the models’ accuracy in generating API calls.
-Notably, the fine-tuned CodeLlama-13b model with API Pack surpasses the performance of GPT-3.5 and GPT-4 models (not fine-tuned) for Level 3, both 0-shot and 3-shot (retre).
-Table 4: Performance comparison for different retrieval methods in three-shot prompting.
+5.2 Retrieval Augmentation Improves API Call Generation
+Table 4 illustrates the impact of different retrieval methods on 3-shot API call generation.
+The results demonstrate that at inference time, 3-shot (retre) is the most effective approach for fine-tuned Mistral-7b and CodeLlama-13b models compared to 3-shot (rand) and 3-shot (retre & rerank).
+These models were fine-tuned using 20,000 Python instances with the 3-shot-tuning template.
+Table 4: Performance comparison for different retrieval methods in 3-shot prompting. Note that 3-shot (retre & rerank) is short for 3-shot retrieved & re-ranked.
 Model Testing Evaluation Accuracy (%)
@@
 3-shot (retre & rerank) 61.0 52.9 55.1 49.2 55.9 49.3
-5.2 Retrieval Augmentation Improves API Call Generation
-Table 4 illustrates the impact of different retrieval methods on three-shot API call generation, highlighting the significance of retrieval augmentation in improving the models’ performance.
-The results of fine-tuning Mistral-7b and CodeLlama-13b models with 20,000 Python instances from API Pack using the three-shot template demonstrate that three-shot retrieval is the most effective approach.
 5.3 Cross-lingual Generalizations in API Call Performance
+Figure 2 compares three fine-tuning approaches for cross-lingual API call performance: a model fine-tuned exclusively on 100,000 instances of ‘cURL’ data, or cURL Model, three expert models, each fine-tuned on 100,000 samples of ‘cURL’, ‘Python’, and ‘Java’ data separately, and a mixture model fine-tuned on 100,000 instances of ‘cURL’ data with additional samples of 1,000 instances each for nine different languages.
+The graph shows that generalizing to new programming languages without fine-tuning is challenging in 0-shot scenarios.
+However, the models perform better than expected in 3-shot settings, indicating some in-context learning adaptability.
+Mixed-language fine-tuning improves performance in both 0-shot and 3-shot scenarios, suggesting that even a small amount of fine-tuning data in various languages can improve the model’s effectiveness in cross-lingual API call tasks.
+These findings demonstrate that cross-lingual API call generation can be achieved by fine-tuning models on a large amount of data in one language plus small amounts of data from other languages, without requiring an extensive dataset for each target programming language.
+Figure 3 illustrates the models’ adaptability across ten programming languages, with the Mixture Model performing on par with Expert Models in their perspective programming languages in 3-shot testing, highlighting their potential in multi-lingual programming applications.
 Figure 2: Cross-lingual API call performance for different models in cURL, Python, and Java.
 Figure 3: Three-shot performance in ten languages.
-Figure 3 compares three fine-tuning approaches for cross-lingual API call performance: a model fine-tuned exclusively on 100,000 instances of ‘cURL’ data, or cURL Model, three expert models, each fine-tuned on 100,000 samples of ‘cURL’, ‘Python’, and ‘Java’ data separately, and a mixture model fine-tuned on 100,000 instances of ‘cURL’ data with additional samples of 1,000 instances each for nine different languages.
-The graph shows that generalizing to new programming languages without fine-tuning is challenging in zero-shot scenarios.
-However, the models perform better than expected in three-shot settings, indicating some in-context learning adaptability.
-Mixed-language fine-tuning improves performance in both zero-shot and three-shot scenarios, suggesting that even a small amount of fine-tuning data in various languages can improve the model’s effectiveness in cross-lingual API call tasks.
-These findings demonstrate that cross-lingual API call generation can be achieved by fine-tuning models on a large amount of data in one language plus small amounts of data from other languages, without requiring an extensive dataset for each target programming language.
-Figure 4 illustrates the models’ adaptability across ten programming languages, with the Mixture Model performing on par with Expert Models in three-shot testing, highlighting their potential in multi-lingual programming applications.
 5.4 Scaling Instruction Dataset Helps Generalization
-Figure 5: Scaling instruction dataset on CodeLlama-13b with zero-shot and three-shot retre evaluations.
+Figure 4: Scaling instruction dataset on CodeLlama-13b with 0-shot and 3-shot retrieval evaluations.
 The x-axis represents the size of fine-tuning data from API Pack.
@@
 Fig.
-5 clearly demonstrates an upward trend in zero-shot performance that correlates with the fine-tuning dataset size (20k, 40k, 80k, 100k, and 1 million instances from API Pack).
-This improvement highlights the advantage of larger datasets in providing diverse examples, which is crucial in zero-shot prompting where the model relies solely on its pre-existing knowledge.
-For three-shot prompting, the graph also shows improvements as the dataset size increases, although less significantly than in the zero-shot case, particular for level 3 cases.
+4 clearly demonstrates an upward trend in 0-shot performance that correlates with the fine-tuning dataset size (20k, 40k, 80k, 100k, and 1 million instances from API Pack).
+This improvement highlights the advantage of larger datasets in providing diverse examples, which is crucial in 0-shot prompting where the model relies solely on its pre-existing knowledge.
+For 3-shot prompting, the graph also shows improvements as the dataset size increases, although less prominent than in the 0-shot case, particularly for level 3 cases.
 This suggests that while additional fine-tuning data is beneficial, model pretraining plays a more critical role in in-context learning ability.
@@
 3-shot 63.5 55.5 56.8 51.4 56.1 49.1
-To verify API Pack’s quality, we evaluated a fine-tuned CodeLlama-13b model on 20k instances of ToolBench dataset, which only contains APIs from RapidAPI.
-Results in Table 5 show that using APIs from a single source like in TollBench decreases model performance in both 0-shot and 3-shot settings compared to API Pack, which curates data from four different sources.
+To demonstrate the benefit of API Pack compared to existing API datasets in the literature, we evaluated a fine-tuned CodeLlama-13b model on 20k instances of ToolBench dataset, which only contains APIs from RapidAPI.
+Results in Table 5 show that using APIs from a single source like in ToolBench significantly decreases model performance in both 0-shot and 3-shot settings compared to API Pack, which curates data from four different sources.
 Table 6: Performance comparison for none-filtered and with-filtered datasets in different shot settings.
@@
 6 Conclusion
-In this paper, we introduce API Pack, a massive multi-programming language dataset with over 1 million instruction-API call instances.
-Using this dataset, we investigate if models can generalize to new languages without extensive multi-language data and whether exposure to diverse APIs during fine-tuning improves generalization.
-Our results show that cross-lingual API call is achievable by training on a large amount of single-language data supplemented with small amounts of data from other languages.
-Furthermore, increasing data volume improves generalization.
+In this paper, we introduced API Pack, a multi-programming language dataset containing over 1 million instruction-API call pairs.
+Our experiments demonstrated that fine-tuning CodeLlama-13B on 20,000 Python instances from API Pack enables it to outperform GPT-3.5 and GPT-4 in generating unseen API calls.
+We also showed that cross-lingual API call generation is achievable by primarily training on single-language data augmented with a small amount of data from other languages.
+Furthermore, we found that increasing the training data volume improves the model’s ability to generalize to unseen API calls.
 These findings highlight the potential of large-scale, diverse datasets in enabling LLMs to effectively identify appropriate API endpoints and generate corresponding API calls based on natural language instructions.
-7 Limitations
-API Pack is not designed for multi-API call scenarios, so models fine-tuned with this dataset may struggle with tasks involving multiple, interdependent API calls.
-This limitation restricts API Pack’s effectiveness in complex, real-world software development tasks that often require multi-API integrations.
-Future work should focus on creating datasets and models that can effectively handle multi-API call scenarios, enabling the generation of more sophisticated and practical code solutions.
+7 Limitations and Future Work
+The first release of API Pack exhibits a set of core limitations.
+First, the fact API Pack was not designed for multi-API call scenarios, limits its application to single-API development tasks.
+Second, current instruction-style requires of the API Name.
+Third, programming language coverage is inherently limited by the library that we used to generate code.
+Finally, only few OAS files contained license information.
+We plan to address these limitations in future versions of the dataset, by integrating multi-API call scenarios and instructions that simply require a task description.
+We also plan to explore other methods to generate API calls so that we can further increase programming language diversity, and to track down complete and accurate license information for all APIs included in API Pack.
 8 Broader Impact Statement